OpenStudy (anonymous):

find the limit of[(3/1-x^3)-(1/1-x)] when x approaches 1

OpenStudy (tkhunny):

By the time you get to Calculus, you should have more of a handle on the Order of Operations. You have just written: \(\left(\dfrac{3}{1} - x^{3}\right) - \left(\dfrac{1}{1} - x\right)\). Is that what you intended or shall we pursue the more interesting problem, \(\dfrac{3}{1-x^{3}} - \dfrac{1}{1-x}\), which you have NOT written?